OpenStudy (anonymous):

Estimate the area under the graph of f(x) = 2 + 2x2 from x = -1 to x = 2 using three rectangles and right end-points

OpenStudy (john_es):

Well, we choose the height the rectangles to be, \[f(x) \Rightarrow x \text{ will be the middle point between the x partition }\] So, our x coordinates will be, \[x_1=(-1+0)/2=-0.5\]\[x_2=(1+0)/2=0.5\]\[x_3=(2+1)/2=1.5\] The base of each rectangle will have 1 unit of longitude, because, \[\text{(End point-First point)/(Number of rectangles)}(2-(-1))/3=1\] So the area will be, \[A\approx1\cdot(f(x_1)+f(x_2)+f(x_3))=11.5\]

OpenStudy (anonymous):

why do you add 0 to -1 and 1 but add a 1 to 2 in order to find the coordinates ?

OpenStudy (john_es):

The idea is to find the middle point of the intervals, where the rectangle is possibly best placed.

OpenStudy (john_es):

You can try other heights: left heights (-1,0,1) or right heights (0,1,2), but these sums are a worse aproximation.

OpenStudy (john_es):

You can check a graphic example here, http://demonstrations.wolfram.com/ComparingBasicNumericalIntegrationMethods/

OpenStudy (anonymous):

Ahh! I see... ok thanks for your help